A shopkeeper marks up his price to make a 30 % profit on cost but allows 5 % discount on cash terms. If he sells an article for $ 247 on cash terms, find his actual profit and express it as a percentage of the cost price of the article.
Answer by rwm(914)

.95*(1.3c)=247
c=200
247-200=$47 profit
47/200=23.5%
[An aside-$50 would be 25% profit]